OpenStudy (anonymous):

A. 2x-10=34 x= B. 3x^3=-24 x= C. (x-9)(x+7)=0 x= D. x^2+2x=24 x=

OpenStudy (jlastino):

A 2x-10=34 2x=34+10 2x=44 x=22 B 3x^3=-24 x^3=-8 x=-2 C (x-9)(x+7)=0 x=9 x=-7 D x^2+2x=24 x^2+2x-24=0 (x-4)(x+6)=0 x=4 x=-6

OpenStudy (anonymous):

Thanks Two right and two wrong. Thanks for showing me how to solve it.

OpenStudy (anonymous):

I mean I had two right and two wrong. Not your answers.